OpenStudy (anonymous):

rewrite 11/12 and 14/15 so that they have a common denominator.

OpenStudy (anonymous):

11/12 = 22/44 and 14/15 = 28/30

OpenStudy (anonymous):

@n1995 it needs to have the same denominator

OpenStudy (anonymous):

The easiest way is to multiply both denominators. If you do that you can just multiply the numerator of the first by the denominator of the second and vice versa. Like that: \[\frac{ a }{ b }+\frac{ c }{ d }=\frac{ a \times d }{ b \times d }+\frac{ c \times b }{ d \times b }=\frac{ ad+cb }{ bd }\]

OpenStudy (anonymous):

what should my answers be?

OpenStudy (anonymous):

Come on, I already gave you de answer, if you understand that you are ok. If I give you the pure answer you will not learn it, try it and post it, if it is wrong I'll tell you.